Zinc reacts with element X to form a compound used as rat poison. In a chemical analysis, it is found that 3.90g zinc combines with 1.24g X to form the compound with empirical formula Zn3X2. Determine the relative atomic mass of element X.


[Relative atomic mass: Zn, 65]

1
Expert's answer
2022-01-10T18:45:03-0500

Solution:

Calculate the moles of Zn:

The molar mass of Zn is 65 g/mol

Therefore,

(3.90 g Zn) × (1 mol Zn / 65 g Zn) = 0.06 mol Zn


Balanced chemical equation:

3Zn + 2X → Zn3X2

According to stoichiometry:

3 moles of Zn reacts with 2 moles of X

Thus 0.06 moles of Zn reacts with:

(0.06 mol Zn) × (2 mol X / 3 mol Zn) = 0.04 mol X


Calculate the relative atomic mass of element X:

(1.24 g X) / (0.04 mol X) = 31 g/mol - phosphorus (P)

Zn3X2 = Zn3P2

The relative atomic mass of element X is 31


Answer: The relative atomic mass of element X is 31